the line y=x is a tangent at (0,0)to a circle with radius 1 , the its centre is???

Grade:12

2 Answers

Arun
6 years ago
Dear student
 
 Let the equation of circle is
(x-a)² + (y-b)² = 1
 
Now circle pass throgh (0,0)
 
a² + b² = 1
 
And also y= x is tangent 
Then one new relationship will also be formed.
 
Hence now you would be able to solve.
Dilli PRASAD
6 years ago
Y=x .....(1) let Centre s of circle be A and B now AB parpendicular to (1) AB slope =-1 Inclination= 135Now A, B are two points which are at a distance 1 (radius) from (0,0) Now A,B=(0+1cos135,0+1sin135) ,(0-cos135,0-sin135)